OpenStudy (anonymous):

What is the difference between f(x)=3x+2 g(x)=x^2 when i want to solve it using these equations: f(g(x)) and g(f(x)) ?

OpenStudy (campbell_st):

f(g(x)) = 3(x^2) + 2 = 3x^2 + 2 g(f(x) = (3x+2)^2 = 9x^2 +12x + 4

OpenStudy (anonymous):

f(g(x)) = 3(x^2) + 2 = 3x^2 + 2 g(f(x) = (3x+2)^2 = 9x^2 +12x + 4

OpenStudy (anonymous):

thanks thats what i got too but i wasnt so sure

OpenStudy (anonymous):

welcome